# ReviewEval: An Evaluation Framework for AI-Generated Reviews

### v2

@@ 3.3 Constructiveness
We assess review constructiveness by quantifying the presence and quality of actionable insights in AI-generated reviews relative to expert feedback. Our framework begins by extracting key actionable components from each review using an LLM with few-shot examples. Specifically, we identify the following actionable insights: (i) criticism points ($C$), which capture highlighted flaws or shortcomings in the paper’s content, clarity, novelty, and execution; (ii) methodological feedback ($M$), which encompasses detailed analysis of experimental design, techniques, and suggestions for methodological improvements; and (iii) suggestions for improvement ($I$), which consist of broader recommendations for enhancement such as additional experiments, alternative methodologies, or improved clarity.
Once these components are extracted, each insight is evaluated along three dimensions: specificity, feasibility, and implementation details. The specificity score $\sigma$ is defined as 1 if the insight is specific and includes explicit examples, and 0 otherwise; the feasibility score $\phi$ is set to 1 when the recommendation is practical within the paper’s context, and 0 otherwise; and the implementation details score $\zeta$ is 1 if actionable steps or detailed methodologies are provided, and 0 otherwise. The overall actionability score for an individual insight is then computed as $S_{\text{act},i}=\sigma_{i}+\phi_{i}+\zeta_{i},$ with an insight considered actionable if $S_{\text{act},i}>0$. Finally, we quantify the overall constructiveness of a review by calculating the percentage of actionable insights:
$S_{\text{act}}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\big{(}S_{\text{act},i}>0\big{)}\times 100,$ (4)
where $N$ is the total number of extracted insights and $\mathbb{I}(\cdot)$ denotes the indicator function. This metric provides a quantitative measure of how effectively a review offers concrete guidance for improving the work.
@@ 3.4 Depth of Analysis
To assess whether a review provides a comprehensive, critical evaluation rather than a superficial commentary, we measure the depth of analysis in AI-generated reviews. This metric captures how thoroughly a review engages with key aspects of a paper, including comparisons with existing literature, identification of logical gaps, methodological scrutiny, interpretation of results, and evaluation of theoretical contributions.
Each review is evaluated by multiple LLMs, which assign scores for each of the five dimensions, $m_{i}$ ($i\in\{1,2,3,4,5\}$), with scores $S_{i}\in[0,1]$. Scores in the continuous range allow us to capture nuances in performance. We define the metrics as follows:
@@
$S_{1}=\begin{cases}3,&\text{if the review provides a thorough, critical comparison,}\\ 2,&\text{if the comparison is meaningful yet shallow,}\\ 1,&\text{if the comparison is vague or lacks specific references,}\\ 0,&\text{if no comparison is provided.}\end{cases}$
@@
$S_{2}=\begin{cases}3,&\text{if the review identifies comprehensive gaps and provides suggestions,}\\ 2,&\text{if it notes some gaps with unclear recommendations,}\\ 1,&\text{if the gaps are vaguely mentioned without solutions,}\\ 0,&\text{if no gaps are identified.}\end{cases}$
@@
$S_{3}=\begin{cases}3,&\text{if the review delivers a thorough critique with actionable suggestions,}\\ 2,&\text{if the critique is meaningful but lacks depth,}\\ 1,&\text{if the critique is vague and offers little insight,}\\ 0,&\text{if no methodological critique is provided.}\end{cases}$
Results interpretation ($m_{4}$): Assesses how well the review interprets the results, addressing biases, alternative explanations, and broader implications:
$S_{4}=\begin{cases}3,&\text{if the interpretation is detailed and insightful,}\\ 2,&\text{if it is meaningful yet shallow,}\\ 1,&\text{if the discussion is generic or vague,}\\ 0,&\text{if no interpretation is offered.}\end{cases}$
@@
$S_{5}=\begin{cases}3,&\text{if the evaluation is comprehensive and insightful,}\\ 2,&\text{if the evaluation is meaningful but lacks depth,}\\ 1,&\text{if the critique is vague,}\\ 0,&\text{if no theoretical assessment is provided.}\end{cases}$
The overall depth of analysis score for a review is calculated as the average normalized score across all dimensions:
$S_{\text{depth}}=\frac{\sum_{i=1}^{5}S_{i}}{15}.$ (5)
@@ 3.5 Adherence to Reviewer Guidelines
To assess whether a review complies with established criteria, we evaluate its adherence to guidelines set by the venue. This metric measures how well the reviewer applies key aspects such as originality, methodology, results, clarity, and ethical considerations, thereby ensuring an objective and structured evaluation process.
@@
$S_{i}=\begin{cases}3,&\text{if there is strong adherence with detailed, accurate feedback;}\\ 2,&\text{if the review shows reasonable alignment with minor deviations;}\\ 1,&\text{if the feedback is incomplete or inaccurate;}\\ 0,&\text{if there is no alignment.}\end{cases}$
@@
$S_{i}=\begin{cases}3,&\text{if the review adheres to the required scale and structure;}\\ 0,&\text{otherwise.}\end{cases}$
The overall adherence score is then computed as:
$S_{\text{adherence}}=\frac{\sum_{i=1}^{2}S_{i}}{6},$ (6)
This normalized score provides a quantitative measure of how well the review conforms to the prescribed guidelines.
@@ 3.6 Conference-Specific Review Alignment
To tailor the review process to conference-specific guidelines, we first retrieve the relevant textual content from the target conference’s official reviewing website using the Extractor API. The extracted text is then processed using GPT Achiam et al. (2023) to filter out extraneous details while preserving essential reviewing instructions. Each guideline $g_{i}$ is converted into a step-by-step instructional prompt via GPT:
$P_{i}=\text{GeneratePrompt}(g_{i}),$
where $P_{i}$ denotes the prompt corresponding to guideline $g_{i}$. Since some criteria apply to multiple sections of a research paper, the prompts are dynamically mapped to the relevant sections. This is achieved using a mapping function:
$S_{j}=\mathcal{M}(P_{i}),$
where $S_{j}$ is the set of paper sections associated with prompt $P_{i}$. Notably, $\mathcal{M}$ is a one-to-many mapping, i.e.,
$\mathcal{M}:P\to\mathcal{P}(S),$
with $\mathcal{P}(S)$ denoting the power set of all sections. By conducting reviews on a section-wise basis, our framework enhances processing efficiency and allows for independent evaluation of each section prior to aggregating the final review. Detailed descriptions of all prompts are provided in the Appendix.
3.7 Review Prompt Iterative Refinement
To enhance the quality and completeness of our review prompts, we employ an iterative refinement process using a Supervisor LLM. Starting with an initial set of prompts generated from the reviewing guidelines, the Supervisor LLM evaluates each prompt in conjunction with its corresponding guideline, serving as the problem statement, and provides targeted feedback. This feedback addresses key aspects such as clarity (ensuring precise and unambiguous instructions), logical consistency (ensuring coherent reasoning), alignment with the guideline, and comprehensiveness (ensuring all relevant aspects are covered). The feedback is then used to revise the prompt, and this iterative loop is repeated for a fixed number of iterations (three in our implementation). The result is a final set of structured, high-quality review prompts that are well-aligned with the reviewing guidelines and optimized for the evaluation process.
Model Actionable Adherence to Coverage Semantic Depth of Factual
@@ 3.7 Review Prompt Iterative Refinement
Sakana-GPT-4o $0.7909\pm 0.1193$ $0.6327\pm 0.0108$ $0.6863\pm 0.1445$ $\mathbf{0.8440\pm 0.0226}$ $0.5208\pm 0.0602$ $0.61\pm 0.014$
Sakana-GPT-4o-Mini $0.7916\pm 0.1246$ $0.6013\pm 0.0669$ $0.7342\pm 0.1799$ $0.8344\pm 0.0161$ $0.4944\pm 0.0402$ $0.60\pm 0.133$
Sakana-3.5-Sonnet $0.6956\pm 0.1475$ $0.6400\pm 0.0297$ $0.7159\pm 0.1407$ $0.8195\pm 0.0169$ $0.4875\pm 0.0737$ $0.72\pm 0.024$
Sakana-3.5-Haiku $0.4634\pm 0.1380$ $0.6313\pm 0.0398$ $0.6641\pm 0.1467$ $0.8130\pm 0.0177$ $0.4597\pm 0.0611$ $0.76\pm 0.0244$
MARG-GPT-4o $0.7253\pm 0.1638$ $0.3206\pm 0.0364$ $0.6978\pm 0.1549$ $0.8005\pm 0.0206$ $\mathbf{0.6833\pm 0.0725}$ $0.73\pm 0.038$
MARG-GPT-4o-mini $0.7428\pm 0.1815$ $0.3031\pm 0.0396$ $0.5439\pm 0.1636$ $0.7926\pm 0.0246$ $0.7014\pm 0.0832$ $0.63\pm 0.084$
Ours-GPT-4o $0.7483\pm 0.1275$ $0.6823\pm 0.0917$ $0.6788\pm 0.2546$ $0.8171\pm 0.0724$ $0.5028\pm 0.0638$ $0.70\pm 0.0457$
Ours-GPT-4o-Mini $0.7385\pm 0.0952$ $\mathbf{0.5785\pm 0.2174}$ $0.6924\pm 0.1650$ $0.8170\pm 0.0346$ $0.5243\pm 0.0628$ $0.68\pm 0.135$
Ours-3.5-Sonnet $\mathbf{0.7513\pm 0.1669}$ $0.6658\pm 0.1180$ $0.6782\pm 0.1468$ $0.8379\pm 0.0204$ $0.5417\pm 0.0918$ $0.65\pm 0.089$
Ours-3.5-Haiku $0.4943\pm 0.2070$ $0.4631\pm 0.1197$ $\mathbf{0.8013\pm 0.1572}$ $0.8274\pm 0.0210$ $0.5597\pm 0.1107$ $\mathbf{0.78\pm 0.0387}$
Expert $0.7522\pm 0.1206$ $0.5708\pm 0.1133$ NA NA $0.6264\pm 0.0722$ $0.90\pm 0.078$
@@
Table 2: Evaluation of AI-generated reviews across six different metrics. Our proposed method is compared against MARG D’Arcy et al. (2024) and Sakana AI Scientist Lu et al. (2024). For each metric, the value closest to the Expert rating is considered best, and such results are highlighted in bold. The ’NA’ for Expert indicates that the metric is calculated relative to the Expert rating, and therefore does not require a comparison

### v3

@@ 3.3 Constructiveness
We assess review constructiveness by quantifying the presence and quality of actionable insights in AI-generated reviews relative to expert feedback. Our framework begins by extracting key actionable components from each review using an LLM with few-shot examples. Specifically, we identify the following insights: (i) criticism points ($CP$), which capture highlighted flaws or shortcomings in the paper’s content, clarity, novelty, and execution; (ii) methodological feedback ($MF$), which encompasses detailed analysis of experimental design, techniques, and suggestions for methodological improvements; and (iii) suggestions for improvement ($SI$), which consist of broader recommendations for enhancement such as additional experiments, alternative methodologies, or improved clarity.
Once these components are extracted, each insight is evaluated along three dimensions: specificity, feasibility, and implementation details.
  • The specificity score $\sigma$ is defined as 1 if the insight is clear and unambiguous, referring to a particular aspect of the paper and including explicit examples.
  • The feasibility score $\phi$ = 1 when the recommendation can realistically be implemented within the research context, such as available data, reasonable technical effort, or domain constraints.
  • The implementation detail score $\delta$ = 1 if the feedback provides actionable steps, specific techniques, or references for improvement.
Examples explaining the three dimensions in Appendix A.3.
The overall actionability score for an individual insight is then computed as $S_{\text{act},i}=\sigma_{i}+\phi_{i}+\zeta_{i},$ with an insight considered actionable if $S_{\text{act},i}>1$ (having atleast 2 of the three qualities mentioned above). Finally, we quantify the overall constructiveness of a review by calculating the percentage of actionable insights:
$S_{\text{act}}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{I}\big{(}S_{\text{act},i}>1\big{)}\times 100,$ (4)
where $N$ is the total number of extracted insights and $\mathbb{I}(\cdot)$ denotes the indicator function. This metric provides a quantitative measure of how a review offers guidance for improving the work.
@@ 3.4 Depth of Analysis
To assess whether a review provides a comprehensive, critical evaluation rather than a superficial commentary, we measure the depth of analysis in AI-generated reviews. Each review is evaluated by multiple LLMs, which assign scores for each dimension, $m_{i}$ ($i\in\{1,2,3,4,5\}$), with scores $S_{i}\in\{0,1,2,3\}$. We define these as follows:
@@
[…]
@@
[…]
@@
[…]
Results interpretation $(m_{4})$: Assesses depth of result discussion (biases, alternative explanations, implications).
[…]
@@
[…]
The overall depth of analysis score for a review is calculated as the average normalized score across all dimensions: $S_{\text{depth}}=\displaystyle\frac{\sum_{i=1}^{5}S_{i}}{15}$
@@ 3.5 Adherence to Reviewer Guidelines
To assess whether a review complies with established criteria, we evaluate its adherence to guidelines set by the venue.
@@
[…]
@@
$S_{i}=\{3\!:\text{ adheres to scale and structure};\;0\!:\text{ otherwise}\}$
The overall adherence score is then computed as $S_{\text{adherence}}=\frac{\sum_{i=1}^{2}S_{i}}{6}$.
This score quantifies how well the review adheres to the prescribed guidelines.
4 ReviewAgent
To ensure AI-based peer reviews are thorough across the multi-dimensional metrics defined in ReviewEval, we introduce ReviewAgent, a framework that aligns evaluations with conference-specific reviewing guidelines while operating effectively across those dimensions.
@@ 4.1 Conference-Specific Review Alignment
To tailor the review process to conference-specific guidelines, we first retrieve them from the target conference’s official reviewing website and pre-process them using LLM. Each review guideline $g_{i}$ is then converted into a step-by-step instructional prompt via LLM: $P_{i}=\text{GeneratePrompt}(g_{i})$, where $P_{i}$ denotes the prompt corresponding to guideline $g_{i}$. Since some criteria apply to multiple sections of a research paper, the prompts are dynamically mapped to the relevant sections. This is achieved using a mapping function: $S_{j}=\mathcal{M}(P_{i})$, where $S_{j}$ is the set of paper sections associated with prompt $P_{i}$. Notably, $\mathcal{M}$ is a one-to-many mapping, i.e., $\mathcal{M}:P\to\mathcal{P}(S)$, with $\mathcal{P}(S)$ denoting the power set of all sections. By conducting reviews on a section-wise basis, our framework enhances processing efficiency and allows for independent evaluation of each section prior to aggregating the final review. (Prompts provided in Appendix A.4)
4.2 Iterative Refinement Loop
To enhance the quality and completeness of our review prompts and the section-wise reviews generated, we employ an iterative refinement process using a Supervisor LLM. Starting with an initial set of prompts generated from the reviewing guidelines, the Supervisor LLM evaluates each prompt in conjunction with its corresponding guideline, serving as the problem statement, and provides targeted feedback. This feedback addresses key aspects such as clarity logical consistency alignment with the guideline, and comprehensiveness. The feedback is then used to revise the prompt, and this iterative loop is repeated for a fixed number of iterations (one in our current implementation). The result is a final set of structured, high-quality review prompts that are well-aligned with the reviewing guidelines and optimized for the evaluation process. We employ the same mechanism to iteratively refine the section-wise reviews generated, ensuring each section benefits from targeted feedback and improvement.
4.3 Improvement using ReviewEval
To further improve the quality and consistency of our reviews, we use an automated improvement loop based on an ReviewEval. After a draft for review is generated for a paper, ReviewEval evaluates it across four important dimensions: (1) constructiveness,(2) depth of analysis, (3) factual accuracy, and (4) adherence to reviewer guidelines a score along with scores of each criteria in these metrics. These are then passed on to the agent, which consumes both the raw review data and the evaluation feedback to generate an enhanced, final review. In our current implementation this loop is run once.
Notably, throughout system development and benchmarking we do conduct an AI vs Human analysis in ReviewEval to judge AI generated reviews against expert reviews taking them as gold standard. But for real world submissions where we do not have any existing reviewer comments, we solely depend on these four objective criteria.
Framework Actionable Adherence to Coverage Semantic Depth of Factual
@@ 4.3 Improvement using ReviewEval
Expert $0.4457\pm 0.1322$ $0.5967\pm 0.1369$ N/A $\pm$ N/A N/A $\pm$ N/A $0.8772\pm 0.1127$ $0.9797\pm 0.0400$
MARG-V1-Deepseek $0.4915\pm 0.1193$ $0.4367\pm 0.1441$ $0.5315\pm 0.2139$ $0.7957\pm 0.0710$ $0.8022\pm 0.1231$ $0.9873\pm 0.0283$
MARG-V1-Qwen $\mathbf{0.6161\pm 0.1098}$ $0.5160\pm 0.1411$ $0.5123\pm 0.2057$ $0.7809\pm 0.0818$ $0.9098\pm 0.0929$ $0.9777\pm 0.0388$
MARG-V2-Deepseek $0.3702\pm 0.1603$ $0.5898\pm 0.1376$ $0.8694\pm 0.1625$ $0.7638\pm 0.0918$ $0.9122\pm 0.0629$ $\mathbf{0.9911\pm 0.0305}$
MARG-V2-Qwen $0.4874\pm 0.1441$ $0.6337\pm 0.1418$ $0.8341\pm 0.1856$ $0.7814\pm 0.0809$ $0.9429\pm 0.0544$ $0.9762\pm 0.0565$
ReviewAgent-Deepseek $0.3096\pm 0.1438$ $0.7584\pm 0.1266$ $0.7330\pm 0.1709$ $0.8016\pm 0.0990$ $0.8694\pm 0.1239$ $0.9870\pm 0.0556$
ReviewAgent-Qwen $0.5899\pm 0.1278$ $\mathbf{0.8545\pm 0.1039}$ $0.7351\pm 0.1803$ $0.7887\pm 0.1089$ $\mathbf{0.9733\pm 0.0514}$ $0.9793\pm 0.0436$
AI-Scientist-Deepseek $0.3164\pm 0.1611$ $0.5997\pm 0.1077$ $\mathbf{0.8913\pm 0.1450}$ $0.7876\pm 0.1012$ $0.7045\pm 0.1320$ $0.9743\pm 0.0966$
AI-Scientist-Qwen $0.4684\pm 0.1641$ $0.7105\pm 0.1508$ $0.7933\pm 0.1696$ $\mathbf{0.8088\pm 0.0988}$ $0.9010\pm 0.0965$ $0.9901\pm 0.0330$
@@
Table 2: Evaluation of AI-generated reviews across six different metrics on FullCorpus-120 dataset. ReviewAgent is compared against MARG D’Arcy et al. (2024) and Sakana AI Scientist Lu et al. (2024). The ’NA’ for Expert indicates that the metric is calculated relative to the Expert rating, and therefore does not require a comparison
Framework Actionable Adherence to Coverage Semantic Depth of Factual
MARG-V1-3.5-Haiku $0.4262\pm 0.1287$ $0.4587\pm 0.1004$ $0.5096\pm 0.1985$ $0.4330\pm 0.0684$ $0.4689\pm 0.1551$ $0.9650\pm 0.0711$
MARG-V1-3.7-Sonnet $0.4725\pm 0.1017$ $0.5820\pm 0.0854$ $0.4355\pm 0.2203$ $0.5057\pm 0.0782$ $0.8244\pm 0.0884$ $0.9593\pm 0.0673$
MARG-V1-GPT4o $0.3762\pm 0.1116$ $0.5567\pm 0.1159$ $0.4870\pm 0.1906$ $0.5019\pm 0.0705$ $0.7178\pm 0.1283$ $0.9587\pm 0.1089$
MARG-V1-GPT4o-mini $0.4338\pm 0.1139$ $0.5433\pm 0.0807$ $0.5034\pm 0.1684$ $0.4974\pm 0.0700$ $0.7867\pm 0.0882$ $0.9877\pm 0.0344$
MARG-V2-3.5-Haiku $0.2746\pm 0.1394$ $0.4400\pm 0.1400$ $\bf{0.9467\pm 0.1383}$ $0.7025\pm 0.1100$ $0.7311\pm 0.0934$ $0.9713\pm 0.0801$
MARG-V2-3.7-Sonnet $0.3384\pm 0.1451$ $0.5047\pm 0.1573$ $0.8817\pm 0.1842$ $0.7639\pm 0.0476$ $0.8044\pm 0.1333$ $0.9903\pm 0.0300$
MARG-V2-GPT4o $0.3077\pm 0.1400$ $0.4957\pm 0.1345$ $0.8896\pm 0.1755$ $0.7763\pm 0.0714$ $0.8156\pm 0.0970$ $0.9900\pm 0.0403$
MARG-V2-GPT4o-mini $0.2137\pm 0.1523$ $0.5167\pm 0.1397$ $0.9361\pm 0.1241$ $0.7701\pm 0.0601$ $0.8644\pm 0.0643$ $0.9887\pm 0.0431$
ReviewAgent-3.5-haiku $0.1781\pm 0.1096$ $0.6867\pm 0.1161$ $0.6868\pm 0.1825$ $0.8380\pm 0.0310$ $0.6911\pm 0.1292$ $\bf{0.9960\pm 0.0219}$
ReviewAgent-3.7-sonnet $0.4135\pm 0.1628$ $0.7213\pm 0.1048$ $0.7915\pm 0.2026$ $0.8202\pm 0.0751$ $0.9222\pm 0.0823$ $0.9837\pm 0.0633$
ReviewAgent-Deepseek-Imp $0.2830\pm 0.1542$ $0.7313\pm 0.1205$ $0.7205\pm 0.1909$ $0.7778\pm 0.1138$ $0.9356\pm 0.0866$ $0.9890\pm 0.0474$
ReviewAgent-Deepseek-ImpRef $0.2927\pm 0.1253$ $0.7330\pm 0.1213$ $0.7369\pm 0.1863$ $0.8021\pm 0.0676$ $0.9533\pm 0.0610$ $0.9943\pm 0.0310$
ReviewAgent-Deepseek-Ref $0.2883\pm 0.1192$ $0.7587\pm 0.1443$ $0.7377\pm 0.1573$ $0.7767\pm 0.1057$ $0.9000\pm 0.0971$ $0.9833\pm 0.0582$
ReviewAgent-GPT4o $0.2740\pm 0.1592$ $0.6773\pm 0.1228$ $0.7565\pm 0.1693$ $0.8018\pm 0.0954$ $0.6889\pm 0.1360$ $0.9833\pm 0.0913$
ReviewAgent-GPT4o-mini $0.2029\pm 0.1391$ $0.6287\pm 0.1235$ $0.7421\pm 0.1600$ $0.8092\pm 0.0906$ $0.6822\pm 0.0954$ $0.9850\pm 0.0634$
ReviewAgent-Qwen-Imp $0.6469\pm 0.1415$ $\bf{0.8787\pm 0.0990}$ $0.6649\pm 0.1727$ $\bf{0.8401\pm 0.0347}$ $0.9822\pm 0.0300$ $0.9707\pm 0.0474$
ReviewAgent-Qwen-ImpRef $\bf{0.6579\pm 0.1194}$ $0.8493\pm 0.1242$ $0.7418\pm 0.1848$ $0.7965\pm 0.0557$ $\bf{0.9889\pm 0.0253}$ $0.9770\pm 0.0341$
ReviewAgent-Qwen-Ref $0.6135\pm 0.1394$ $0.8347\pm 0.1058$ $0.6605\pm 0.1861$ $0.8185\pm 0.0952$ $0.9778\pm 0.0535$ $0.9590\pm 0.0471$
AI-Scientist-3.5-Haiku $0.3257\pm 0.1448$ $0.5360\pm 0.0941$ $0.8624\pm 0.1441$ $0.7786\pm 0.0696$ $0.6133\pm 0.1071$ $0.9953\pm 0.0256$
AI-Scientist-3.7-Sonnet $0.3215\pm 0.1439$ $0.7980\pm 0.1071$ $0.8208\pm 0.1694$ $0.7986\pm 0.0978$ $0.9511\pm 0.0741$ $0.9667\pm 0.1269$
AI-Scientist-GPT4o $0.3071\pm 0.1471$ $0.5217\pm 0.1100$ $0.9079\pm 0.1282$ $0.7661\pm 0.0818$ $0.5911\pm 0.1144$ $0.9640\pm 0.1148$
AI-Scientist-GPT4o-mini $0.3406\pm 0.1390$ $0.5107\pm 0.0957$ $0.8279\pm 0.1776$ $0.7947\pm 0.0417$ $0.6111\pm 0.0944$ $0.9557\pm 0.1903$
Table 3: Evaluation of AI-generated reviews across six different metrics on CoreCorpus-30 dataset. ReviewAgent against MARG and Sakana AI Scientist. ReviewEval framework has two binary settings: Imp (improvement decribed in 4.3) and Ref (reflection described in 4.2). A suffix of -Imp, -Ref or -ImpRef indicates that the corresponding feature was turned “on”; absence of any suffix implies both features were “off.”